OpenStudy (anonymous):

Sarah has a total of 52 DVD's and CD's. If the number of CD's is two more than four times the number of DVD's, how many CD's does she have?

OpenStudy (anonymous):

assume the number of dvd= x then the number of cd =4x+2

OpenStudy (anonymous):

and accordong to the qus total no of cd + dvd=52

OpenStudy (anonymous):

LOL

OpenStudy (anonymous):

so x+4x+2=52

OpenStudy (anonymous):

opps

OpenStudy (anonymous):

now solve it for x

OpenStudy (anonymous):

@LRGAUS95 are u getting...?

OpenStudy (anonymous):

\[5x+2=52\] \[5x=50\] \[x=10\] finish it :)

OpenStudy (anonymous):

using @anuragtripathi 's equation

OpenStudy (anonymous):

thats what i got @ConradBarczyk

OpenStudy (anonymous):

wait its not 5x its 4x

OpenStudy (anonymous):

x=10 is the value of dvd and no of cd as we assumed above =4x+2 put here the value of x=10 so no of cd she have=4x+2 =4*10 +2 =42

OpenStudy (anonymous):

so she has 42 cds @anuragtripathi ?

OpenStudy (anonymous):

yeah correct 42 cds

OpenStudy (anonymous):

thankyou @anuragtripathi

OpenStudy (anonymous):

welcome...:)